OpenStudy (anonymous):

What is the simplified form of the expression:

OpenStudy (anonymous):

\[6\sqrt{v} - \sqrt{4v^2} - \sqrt{36v} + \sqrt{v^2}\]

OpenStudy (bahrom7893):

@dumbcow, go for it :)

OpenStudy (anonymous):

@bahrom7893

OpenStudy (bahrom7893):

Well first of all, I would rewrite this a little: \[6\sqrt{v}-\sqrt{2^2*v^2}-\sqrt{6^2*v}+\sqrt{v^2}\]and keep in mind that\[\sqrt{a^2}=a\]

OpenStudy (bahrom7893):

So after we simplify the top equation a bit, we get: \[6\sqrt{v}-2v-6\sqrt{v}+v=-2v+v=-v\]

OpenStudy (anonymous):

Thank you! I almost got confused for a minute after I read that. Then I got it :P